Please, please help!
After you complete this exam, you and two friends decide to go to McDonald’s to celebrate. Last month McDonald’s filled 90% of the orders accurately. What is the probability that:
a) All three orders will be filled accurately?
b) None of the three orders will be filled accurately?
c) At least 2 of the 3 orders will be filled accurately?

Answer by stanbon(48568)

Last month McDonald’s filled 90% of the orders accurately. What is the probability that:
a) All three orders will be filled accurately?
Prob(3 correct)=0.9^3=0.729
b) None of the three orders will be filled accurately?
P(3 incorrec)= 0.1^3=0.001
c) At least 2 of the 3 orders will be filled accurately?
P(2 or 3 correct)=3C2(0.9)^2(0.1) + 3C3(0.9^3
=3*0.81*0.1+13(0.0729)
=0.243+0.729
=0.972
Cheers,
Stan H.